OpenStudy (anonymous):

If we flip 6 coins, what is the probability of getting at least 2 heads?

OpenStudy (kc_kennylau):

It is 1 - the probability of getting one head or no head

OpenStudy (kc_kennylau):

\[P(X\ge2)\\=1-P(X<2)\\=1-P(X=0)-P(X=1)\]

OpenStudy (anonymous):

so 1-1/2+0=1/2?

OpenStudy (kc_kennylau):

nope

OpenStudy (anonymous):

why not

OpenStudy (kc_kennylau):

(a) What is the probability of getting exactly no head in 6 flips? (b) What is the probability of getting exactly one head in 6 flips?

OpenStudy (anonymous):

for a it would be 1/2^6

OpenStudy (anonymous):

right?

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

and for b i don;t rlly know

OpenStudy (kc_kennylau):

How many cases match (b)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

so 1/2^6?

OpenStudy (kc_kennylau):

Can you list out all the cases that have exactly one head?

OpenStudy (kc_kennylau):

Using H to represent head and T for tail

OpenStudy (anonymous):

wait 6

OpenStudy (anonymous):

6/2^6

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

and do i add them up?

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

so 7/64

OpenStudy (anonymous):

and then i subtract it from 1 right?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

thx

OpenStudy (kc_kennylau):

Full solution: \[P(X\ge2)\\=1-P(X<2)\\=1-P(X=0)-P(X=1)\\=1-\frac1{2^6}-\frac6{2^6}\\=1-\frac7{2^6}\\=\frac{57}{64}\]

OpenStudy (anonymous):

also can u help me on another problem

OpenStudy (kc_kennylau):

sure, open another problem

OpenStudy (anonymous):

Randy presses RAND on his calculator twice to obtain two random numbers between 0 and 1. Let $p$ be the probability that these two numbers and 1 form the sides of an obtuse triangle. Find $4p$.

OpenStudy (kc_kennylau):

So, do you know how to check if three numbers form an obtuse triangle?

OpenStudy (anonymous):

i know that the regular triangle is a+b>c b+c>a a+c>b

OpenStudy (anonymous):

is a obtuse triangle different

OpenStudy (kc_kennylau):

a+b>c is for any triangle

OpenStudy (kc_kennylau):

An obtuse triangle is a^2+b^2<c^2

OpenStudy (kc_kennylau):

where a and b are the shorter sides

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so I plugin every number i think is possible??

OpenStudy (kc_kennylau):

I think you need to draw a graph for this

OpenStudy (anonymous):

can u explain this why

OpenStudy (kc_kennylau):

So it is given that \(0<x<1\) and \(0<y<1\)

OpenStudy (kc_kennylau):

And you are trying to find the probability that \(x^2+y^2<1\)

OpenStudy (anonymous):

so 1 would be the boundaries

OpenStudy (anonymous):